OpenStudy (anonymous):

Estimate: twenty-nine forty-firsts times one hundred is approximately equal to what number? twenty-five fifty seventy-five one hundred

jimthompson5910 (jim_thompson5910):

29 is close to 30 41 is close to 40 So 29/41 is close to 30/40 = 3/4 = 0.75 Multiply this by 100 by moving the decimal point two spots to the right to get 75 So the answer is approximately 75